QUESTION 2 1 points Save Answer Bob Katz and Sally Mander are a married couple with four children. Total wages for 2015 equaled $96,400. Stock which had been purchased nine months earlier was sold for a $1,600 gain and stock held for three years was sold for a $7,000 gain. Interest income from savings was $100. ltemized deductions totaled $29,500. Bob and Sally qualify for a $1,500 tax credit. What is Bobs and Sallys taxable income? O $39,000 O $44,600 O $46,000 O $48,400
Answer #1

Gross income means total income by whatever source, Gross income is not limited to cash received, but "includes income realized in any form, whether money, property, or services.Gross income includes wages and tips, fees for performing services, gain from sale of inventory or other property, interest, dividends, rents, royalties, pensions, alimony, and many other types of income items must be included in income when received or accrued. The amount included is the amount the taxpayer is entitled to receive. Gains on property are the gross proceeds less amounts returned, cost of goods sold, or tax basis of property sold.

The standard deduction is a dollar amount that non-itemizers may subtract from their income before income tax is applied. Taxpayers may choose either itemized deductions or the standard deduction, but usually choose whichever results in the lesser amount of tax payable. The standard deduction is available to US citizens and aliens who are resident for tax purposes and who are individuals, married persons, and heads of household. The standard deduction is based on filing status and typically increases each year. It is not available to nonresident aliens residing in the United States.

a personal exemption is an amount that a resident taxpayer is entitled to claim as a tax deduction against personal income in calculating taxable income and consequently federal income tax. In 2017, the personal exemption amount was $4,050, though the exemption is subject to phase-out limitations.

A. Taxable Income = 96400+1600+7000+100=  $105100

B. itemized deeuction = $29500( assumed joint return filed)

C. personal deduction = 4050×4= $16200

Taxable Income = 105100-29500-16200= 59400

Taxable Income =